You are running a t-test for means. You find a test statistic of -2.4. Your sample size is 7.  Your alternative hypothesis is that the population mean is different from 0. What is a possible range of p-values for your test?

between 0.01 and 0.025

between 0.025 and 0.05

between 0.05 and 0.1

between 0.1 and 0.25

Greater than 0.5
Answer #1

Solution.

The null and alternative hypothesis is ,

H0 :   = 0

Ha :     0

This is the left tailed test

Test statistic = t = -2.4

n = 7

degrees of freedom = n - 1 = 7 - 1 = 6

P( t < -2.4 ) = 0.027

P-value = 2 * P( t < -2.4 )

P-value = 2 * 0.027

P-value = 0.054

between 0.05 and 0.1